# Distribution-Level Feature Distancing for Machine Unlearning: Towards a Better Trade-off Between Model Utility and Forgetting

### v3

@@ Abstract
With the explosive growth of deep learning applications and increasing privacy concerns, the right to be forgotten has become a critical requirement in various AI industries. For example, given a facial recognition system, some individuals may wish to remove their personal data that might have been used in the training phase. Unfortunately, deep neural networks sometimes unexpectedly leak personal identities, making this removal challenging. While recent machine unlearning algorithms aim to enable models to forget such data, we observe an unintended utility drop, termed correlation collapse, where these algorithms inadvertently weaken the essential correlations between image features and true labels during the forgetting process. To address this challenge, we propose Distribution-Level Feature Distancing (DLFD), a novel method that efficiently forgets instances while preserving task-relevant feature correlations. Our method synthesizes data samples by optimizing the feature distribution to be distinctly different from that of forget samples, achieving effective results within a single training epoch. Through extensive experiments on facial recognition datasets, we demonstrate that our approach significantly outperforms state-of-the-art machine unlearning methods in both forgetting performance and model utility preservation. The implementation of DLFD can be accessed at https://github.com/Dasol-Choi/DLFD.
@@ Introduction
Deep neural network models have achieved remarkable success in various computer vision applications (He et al. 2016; Tan and Le 2019; Han et al. 2022; Huang et al. 2017; Jiang et al. 2022). Especially, recent works show large-scale foundation models demonstrate superior classification performance across a range of tasks (Radford et al. 2021; Kolesnikov et al. 2020; Floridi and Chiriatti 2020; Han et al. 2022; Liu et al. 2023). However, alongside these advancements, concerns emerge regarding the unintentional leakage of sensitive information, such as personal identities from training data (Shokri et al. 2017a; Tarun et al. 2023; Hu et al. 2022). This issue becomes particularly pressing in the context of training data extraction attacks, where studies show that even large foundation models leak information when deployed in real-world settings (Yu et al. 2023; Carlini et al. 2021).
@@
Machine unlearning has emerged as a promising solution to mitigate potential data leakage (Tarun et al. 2023; Golatkar, Achille, and Soatto 2020b; Gupta et al. 2021; Bourtoule et al. 2021; Foster, Schoepf, and Brintrup 2024), particularly in upholding the right to be forgotten, which allows individuals to request the removal of their personal information from trained models. For example, in medical AI applications, a patient might request that their medical images, used during the training of a diagnostic model, be removed to protect their privacy. In such a scenario, machine unlearning would enable the model to forget the patient’s data without compromising the model’s overall performance on other tasks. This growing need for privacy has driven interest in machine unlearning research within various AI-driven industries.
@@
Despite the progress made by recent machine unlearning algorithms, our work identifies a critical issue that has not been fully explored: the risk of correlation collapse. When simply applying existing error-maximizing methods (Tarun et al. 2023; Kurmanji, Triantafillou, and Triantafillou 2023; Chundawat et al. 2023b), unexpected outcomes can occur. For instance, these methods can inadvertently increase the magnitude of loss excessively, leading to additional data leakage by making certain data points appear special. Moreover, relying solely on these approaches might degrade the generalization performance of the model on the original task, thereby introducing a trade-off between model utility and forgetting. We believe this degradation is due to correlation collapse, where the useful correlations between image features and their true labels are weakened. To prevent these unexpected performance drops, it is crucial to carefully adapt and improve upon the existing methods.
@@
To address this challenge, we propose a novel framework called Distribution-Level Feature Distancing (DLFD) that enables unlearning of specific images while maintaining the accuracy of the original task. Our approach shifts the feature distribution of the retain images away from the distribution of the images to be forgotten, by leveraging the Optimal Transport (OT) problem (Peyré, Cuturi et al. 2019; Le et al. 2021; Cuturi 2013; Altschuler, Niles-Weed, and Rigollet 2017). Specifically, DLFD generates perturbed images by maximizing the distance between the optimized data distribution and the distribution of the data to be forgotten in the feature space using OT loss.
@@
Our method demonstrates superior performance compared to state-of-the-art (SOTA) methods in a problem setting that closely resembles real-world scenarios. Additionally, we introduce and analyze the concept of correlation collapse, which has not been extensively addressed in previous work, and revisit the task-agnostic instance unlearning setting.
Our contributions are as follows:
@@
On the other hand, some methods focus on data manipulation. For example, UNSIR (Tarun et al. 2023) generates artificial noise that is added to the data to maximize the loss values for a specific target class that needs to be forgotten. Training on these error-maximized data points has shown good forgetting performance. Building on this, another method (Chundawat et al. 2023b) uses samples to be retain to improve unlearning scores, extending the work of UNSIR. Similarly, recent works (Cha et al. 2024) use intentionally perturbing noise to increase the loss value, focusing primarily on error-maximizing synthesized images to achieve a high forgetting score.
@@
Figure 1: The conceptual illustration of the correlation collapse. If following the misguided forgetting direction, the correlation between the task-related useful features and labels can weaken.
@@
In the general computer vision domain, a feature vector $w\in\mathcal{W}$ corresponding to an image $x_{\text{retain}}$ might contain various semantic information (Na, Ji, and Kim 2022; Richardson et al. 2021). Some of these semantic features, which we denote as $w_{\text{task}}$, are highly correlated with the original task that the model $\theta_{\text{original}}$ is designed to solve. However, for personal identity unlearning tasks, another set of features, $w_{\text{identity}}$, represents information specific to a particular personal information.
@@
As illustrated in Figure 2, we can observe this correlation collapse through feature visualization. The Original Model (Figure 2(a)) maintains clear class separations in feature space, indicating well-preserved task-relevant features. In contrast, the Error Maximized model (Figure 2(b)) shows clustered features with diminished class distinctions—a direct result of disrupting the shared feature space while attempting to remove identity information.
(a) Original Model Class-Separated Features
(b) Error Maximized Class-Separated Features
Figure 2: Feature representations visualized from age classification model. (a) demonstrates clear class distinctions, with each class (age group) well-separated in feature space. (b), derived using the Negative Gradient method, shows tightly clustered features with less class distinction, illustrating correlation collapse.
To address this challenge, we propose Distribution-Level Feature Distancing (DLFD), which maintains class-separating task-related features during unlearning (Figure 3). Our method carefully preserves the structure of $\mathcal{W}_{\text{task}}$ while modifying identity-related features. Additionally, we introduce classification loss and dynamic forgetting strategies that adapt the forgetting process based on the model’s ongoing performance, ensuring effective unlearning while maintaining discriminative task-relevant features.
@@
Figure 3: The core method of our framework Distribution-Level Feature Distancing (DLFD): feature distribution optimization through optimal transport. This component generates a synthesized dataset by maximizing the distance between retain and forget data distributions in the feature space. When combined with classification loss preservation and dynamic forgetting strategy (detailed in Algorithm1), DLFD achieves an effective balance between model utility and forgetting performance.
Proposed Methods
In this section, we introduce Distribution-Level Feature Distancing (DLFD), our comprehensive framework for effective machine unlearning. DLFD consists of three key components designed to balance forgetting performance and model utility. Below, we detail each component and explain how they work together to achieve our objectives.
Feature Distribution Optimization
Our first component operationalizes the concept of shifting the retain data distribution ($\mu$) away from the forget data distribution ($\nu$) by leveraging the optimal transport (OT) distance. Unlike simpler metrics such as KL divergence or JS divergence, the OT distance captures the complex, high-dimensional relationships between data points (Arjovsky, Chintala, and Bottou 2017; Gulrajani et al. 2017).
The OT distance between the distributions $\mu$ and $\nu$ is formally defined as:
Here, $\gamma$ represents the set of all possible joint distributions that can transport $\mu$ to $\nu$, and $c(w,w^{\prime})$ represents the cost based on cosine similarity between feature vectors. Given the complexity of directly solving this problem, we employ a differentiable Sinkhorn method (Cuturi 2013; Altschuler, Niles-Weed, and Rigollet 2017) to approximate the solution efficiently. This method reduces the computational complexity to approximately quadratic time $\mathcal{O}(n^{2})$, making it feasible for use in mini-batch computations.
In this equation, the cost matrix $C$ captures the pairwise distances between feature vectors from $\mu$ and $\nu$. The regularization term $\lambda$ ensures that the transport plan $T$ remains smooth and avoids overly concentrated mass transfers, which could destabilize the model. By iteratively optimizing $T^{\lambda}$, the feature distribution optimization effectively separates the retain data distribution from the forgotten data, thus mitigating the risk of correlation collapse.
@@
To maintain model utility during the unlearning process, our second component incorporates a classification loss that guides the perturbation process to address correlation collapse. The classification loss ensures that the original class information of the retain data is preserved, even as the model attempts to forget specific data points. A critical aspect of this component is the use of a linear weight that dynamically adjusts the importance of the classification loss throughout the training process.
The linear weight plays a crucial role in balancing the trade-off between maximizing the separation of distributions and preserving the model’s utility. At the beginning of training, it is set lower, allowing the model to focus more on maximizing the distance between the retain and forgot data distributions. As training progresses, the linear weight gradually increases, shifting the model’s focus toward preserving the original class-specific features of the retain data.
The perturbation applied to the retain data points $x_{i}$ is computed as follows:
@@
Here, $l_{OT}$ represents the OT loss between the retain and forget data distributions, while $l_{CE}$ represents the classification loss, which is weighted by the linear factor $\lambda$ and computed as:
@@
Table 1: Overall performance of various machine unlearning methods on ResNet18 (He et al. 2016) classification tasks. Our method achieves superior NoMUS scores across all tasks, with remarkable forgetting scores while maintaining competitive test accuracy. The best score is in boldface except for the ground-truth (Retrained).
Note: Fine-tuning (Golatkar, Achille, and Soatto 2020b), NegGrad (Golatkar, Achille, and Soatto 2020c), CF-k (Goel, Prabhu, and Kumaraguru 2022a), EU-k (Goel, Prabhu, and Kumaraguru 2022a), UNSIR (Tarun et al. 2023), BadTeaching (Chundawat et al. 2023a), SCRUB (Kurmanji, Triantafillou, and Triantafillou 2023)
@@
The goal of an ideal machine unlearning method is to remove the images that need to be forgotten ($\mathcal{D}_{forget}$) while maintaining the original classification performance. In this study, we adopt a common machine unlearning setting where the model has access to a subset of the training data, $\mathcal{D}_{retain}$, which the AI company may still possess. Formally, we assume that the training data $\mathcal{D}_{train}$ is composed of $\mathcal{D}_{retain}$ and $\mathcal{D}_{forget}$, following the general machine unlearning setting described in (Choi and Na 2023).
Our objective is to develop a machine unlearning algorithm that makes the unlearned model $\theta_{unlearned}$ as similar as possible to the retrained model $\theta_{retrained}$, which is considered the ground truth and is trained only on $\mathcal{D}_{retain}$.
We also introduce a dataset $\mathcal{D}_{unseen}$, which is never used during the training or testing phases of the model. This dataset serves as our test set $\mathcal{D}_{test}$ and is exclusively used for evaluating the forgetting score.
It is important to note that any subject targeted for unlearning should not simultaneously belong to the three datasets: $\mathcal{D}_{forget}$, $\mathcal{D}_{retain}$, and $\mathcal{D}_{unseen}$. This ensures that the subject to be forgotten is not present across multiple datasets in the machine unlearning setting.
@@
In this work, we adopt a task-agnostic machine unlearning setup, where unlearning specific target subjects does not alter the functional scope of the original task. Traditional machine unlearning research has primarily focused on class-unlearning, where entire categories (classes) are removed from the model upon a data removal request (Tarun et al. 2023; Golatkar, Achille, and Soatto 2020b; Goel, Prabhu, and Kumaraguru 2022b). While this approach works in certain scenarios, it is not applicable in all cases. For instance, in a gender classification model, removing the male class would leave only the female class, rendering the model ineffective for its intended purpose of gender classification. Hence, class-unlearning is not always representative of real-world needs.
To address these limitations, we propose an instance-unlearning problem setting, which targets the removal of specific personal identities or data samples without changing the overall function of the model. This approach ensures that the model’s core functionality remains intact, making it more applicable to scenarios where the goal is to forget specific data without compromising the model’s utility (Eleni Triantafillou 2023; Choi and Na 2023; Choi et al. 2024).
@@
Given $\mathcal{D}_{forget}$ and $\mathcal{D}_{unseen}$ datasets, we train a binary classifier $\psi(\cdot)$ to distinguish between them:
@@
  • Age Estimation: We use the MUFAC dataset (Choi and Na 2023), which consists of 13,068 facial images at a resolution of 128$\times$128 pixels. The images are categorized into 8 distinct age groups, and each image is annotated with a unique identity number to facilitate identity-based unlearning. The training set comprises 10,025 samples, with 8,525 retained and 1,500 designated for forgetting, providing a robust basis for evaluating unlearning performance across age-related features.
  • Emotion Recognition: The RAF-DB dataset (Li, Deng, and Du 2017), which we employ for emotion recognition, contains 15,000 facial images. The images are labeled across 7 emotional classes and are characterized by a wide variability in factors such as age, gender, ethnicity, and facial occlusions. The training set consists of 11,044 samples, with 7,730 retained and 3,314 designated for forgetting, making it well-suited for testing the unlearning of emotional expressions under diverse conditions.
  • Multi-Attribute Classification: For multi-attribute classification, we utilize the MUCAC dataset (Choi and Na 2023), derived from CelebA (Liu et al. 2018). This dataset includes 30,000 facial images. Each image is annotated with three binary attributes: gender (female/male), age (old/young), and expression (smiling/unsmiling). The training set comprises 25,933 samples, with 15,385 retained and 10,548 designated for forgetting. This dataset allows for a comprehensive evaluation of unlearning across multiple facial attributes simultaneously. Additionally, for gender classification, we specifically focus on the gender attribute from this dataset as a separate binary classification task.
@@
For experiments, we have utilized various deep-neural network architectures including ResNet (He et al. 2016), DenseNet (Huang et al. 2017), EfficientNet (Tan and Le 2019) which have been largely adopted as the most representative architectures in the common computer vision domain. For a fair comparison, we use the identical $\theta_{original}$ as the starting point for all the machine unlearning methods individually for each task. Specifically, in our experiments, all the machine unlearning methods fine-tune the $\theta_{unlearned}$ that is initially the same as the $\theta_{original}$ in the starting point except the Retrained model. The Retrained model, which serves as the ground truth, is trained from scratch using only the $\mathcal{D}_{retain}$, excluding any data that is intended to be forgotten. This ensures that the retrained model fully represents the desired unlearning outcome.
Given the computational complexity of our proposed DLFD method, which involves calculating OT loss and performing MIA evaluations, we limit the training to a single epoch. Other machine unlearning methods were also trained for 1-2 epochs to ensure a fair comparison. Additionally, we found that a learning rate from 0.001 to 0.005 was effective across all models and methods, a setting consistent with previous work (Tarun et al. 2023).
Performance of DLFD Method
For each deep learning architecture, we train four types of classification models: (1) facial age prediction models, (2) facial emotion recognition models, (3) facial multi-attribute classification models, and (4) gender classification models. The multi-attribute model includes three different binary labels: gender (female/male), age (old/young), and expression (smiling/unsmiling), so we report the average classification accuracy as the model utility. The gender classification model is evaluated as a binary classification task.
Across all experiments, our proposed method consistently delivers competitive or superior performance in both metrics. We demonstrate that DLFD effectively unlearns the samples to be forgotten while maintaining model utility. As illustrated in Figure 5(c), the loss distributions of $\mathcal{D}{unseen}$ and $\mathcal{D}{forget}$ in our proposed DLFD method closely resemble those of the Retrained model, which is considered the ground truth (Figure 5(b)). This high similarity between the loss distributions suggests that the unlearning algorithm has been effectively executed. Moreover, Figure 4 displays the feature representations extracted by the DLFD model. The figure demonstrates how DLFD maintains clear class distinctions, effectively preventing Correlation Collapse and preserving essential task-related features.
In particular, the DLFD method shows more substantial improvements in complex, multi-class tasks such as age estimation and emotion recognition, where the feature entanglement is more significant. Conversely, the improvements in multi-attribute and gender classification tasks are relatively smaller, likely due to the binary nature of these classifications, where the complexity of feature entanglement is inherently lower. These results highlight the effectiveness of our approach in scenarios where maintaining feature integrity amidst complex and overlapping feature spaces is more challenging.
Figure 4: Feature representations from the age classification model using DLFD. The model maintains class separation similar to the original model (Figure 2(a)), showing DLFD’s capability to retain task-relevant features while avoiding the correlation collapse.
@@
Figure 5: The loss distributions for two baselines and ours. In each plot, the orange space represents the loss distribution for unseen data, while the green space represents the loss distribution for forget data. (a) illustrates loss distributions for Original model. (b) shows loss distributions for Retrained model. Finally (c) represents loss distributions for $\theta_{unlearned}$ fine-tuned on our proposed DLFD-based optimized images.
@@
We conducted ablation studies to evaluate the contribution of each component in DLFD. Initially, we assessed the performance using only feature distribution optimization, which achieved moderate success in machine unlearning, as shown in Table 3. While this component is effective in distancing the retain data distribution from the forget data distribution, it can lead to a reduction in model utility if not combined with other components.
Feature Dist. Optimization Classification Loss Dynamic Forgetting Age Emotion Multi-Attr. Gender NoMUS $\uparrow$ NoMUS $\uparrow$ NoMUS $\uparrow$ NoMUS $\uparrow$ 0.7021 0.7199 0.9423 0.8898 0.7536 0.7536 0.9420 0.9170 0.7698 0.7934 0.9433 0.9192
Table 3: Ablation study showing the impact of each component in DLFD. Each row shows the cumulative effect of adding components, with the final row representing the complete DLFD framework.
Next, we incorporate classification loss preservation as the second component, which helps maintain the classification accuracy of the retain data. This addition led to a significant improvement in performance across all tasks, particularly for the Age and Emotion tasks, where the NoMUS score increased by approximately 7.3% (from 0.7021 to 0.7536) and 4.7% (from 0.7199 to 0.7536), respectively. This indicates that classification loss preservation is crucial for maintaining model utility while achieving effective unlearning.
Finally, we integrate dynamic forgetting as the third component, which adapts the unlearning process based on the model’s performance during training. With this addition, we observed further enhancements in performance. For instance, the NoMUS score for the Age and Emotion tasks increased by an additional 2.2% (from 0.7536 to 0.7698) and 5.3% (from 0.7536 to 0.7934), respectively. This component effectively prevents correlation collapse by dynamically adjusting the balance between forgetting and utility preservation.
The results demonstrate that the complete DLFD framework, with all three components together, yields the best overall performance. The synergy between feature distribution optimization, classification loss preservation, and dynamic forgetting enables DLFD to achieve superior NoMUS scores across all tasks compared to using any subset of components.

### v4

@@ Abstract
With the explosive growth of deep learning applications and increasing privacy concerns, the right to be forgotten has become a critical requirement in various AI industries. For example, given a facial recognition system, some individuals may wish to remove their personal data that might have been used in the training phase. Unfortunately, deep neural networks sometimes unexpectedly leak personal identities, making this removal challenging. While recent machine unlearning algorithms aim to enable models to forget specific data, we identify an unintended utility drop—correlation collapse—in which the essential correlations between image features and true labels weaken during the forgetting process. To address this challenge, we propose Distribution-Level Feature Distancing (DLFD), a novel method that efficiently forgets instances while preserving task-relevant feature correlations. Our method synthesizes data samples by optimizing the feature distribution to be distinctly different from that of forget samples, achieving effective results within a single training epoch. Through extensive experiments on facial recognition datasets, we demonstrate that our approach significantly outperforms state-of-the-art machine unlearning methods in both forgetting performance and model utility preservation.
@@ Introduction
Deep neural network models have achieved remarkable success in various computer vision applications (He et al. 2016; Tan and Le 2019; Han et al. 2022; Huang et al. 2017; Jiang et al. 2022). Especially, recent works show large-scale foundation models demonstrate superior classification performance across a range of tasks (Radford et al. 2021; Kolesnikov et al. 2020; Floridi and Chiriatti 2020; Han et al. 2022; Liu et al. 2023). However, alongside these advancements, concerns emerge regarding the unintentional leakage of sensitive information, such as personal identities from training data (Shokri et al. 2017a; Hu et al. 2022).
@@
Machine unlearning has emerged as a promising solution to mitigate potential data leakage (Tarun et al. 2023; Golatkar, Achille, and Soatto 2020b; Gupta et al. 2021; Bourtoule, Chandrasekaran et al. 2021; Foster, Schoepf, and Brintrup 2024), particularly in upholding the right to be forgotten, which allows individuals to request the removal of their personal information from trained models. For example, in medical AI applications, a patient might request that their medical images, used during the training of a diagnostic model, be removed to protect their privacy. In such a scenario, machine unlearning enables the model to forget the patient’s data without compromising overall performance on other tasks. This growing need for privacy has driven interest in machine unlearning research within various AI-driven industries.
@@
Despite advancements in machine unlearning algorithms, we identify a critical issue that has not been fully explored: the risk of correlation collapse. When simply applying existing error-maximizing methods (Tarun et al. 2023; Kurmanji, Triantafillou, and Triantafillou 2023; Chundawat et al. 2023b), unexpected outcomes can occur. For instance, these methods can inadvertently increase the magnitude of loss excessively, leading to additional data leakage by making certain data points appear special. Moreover, relying solely on these approaches may degrade the generalization performance of the model on the original task, introducing a trade-off between model utility and forgetting. We believe this degradation is due to correlation collapse, where the useful correlations between image features and their true labels are weakened. To prevent these unexpected performance drops, it is crucial to carefully adapt and improve upon the existing methods.
@@
To address this challenge, we propose a novel framework, Distribution-Level Feature Distancing (DLFD) that enables unlearning of specific images while maintaining the accuracy of the original task. Our approach shifts the feature distribution of the retain images away from the distribution of the forget images, by leveraging the Optimal Transport (OT) problem (Peyré, Cuturi et al. 2019; Le et al. 2021; Cuturi 2013; Altschuler, Niles-Weed, and Rigollet 2017). Specifically, DLFD generates perturbed images by maximizing the distance between the optimized data distribution and the forget data distribution in the feature space using OT loss.
@@
Our method demonstrates superior performance compared to state-of-the-art methods in a setting that closely reflects real-world scenarios. We also introduce and analyze the concept of correlation collapse, which has not been extensively addressed in previous works, and revisit the task-agnostic instance unlearning setting. Our contributions are as follows:
@@
On the other hand, some methods focus on data manipulation. For example, UNSIR (Tarun et al. 2023) generates noise that is added to the data to maximize the loss values for a specific target class that needs to be forgotten. Training on these error-maximized data points has shown good forgetting performance. Building on this, another method (Chundawat et al. 2023b) uses samples to be retain to improve unlearning scores, extending the work of UNSIR. Similarly, recent works (Cha et al. 2024) use perturbing noise to increase the loss value, focusing primarily on error-maximizing synthesized images to achieve a high forgetting score.
@@
Figure 1: The concept of correlation collapse. If following the misguided forgetting direction, the correlation between the task-related useful features and labels can weaken.
@@
In the general computer vision domain, a feature vector $w\in\mathcal{W}$ corresponding to an image $x$ might contain various semantic information (Na, Ji, and Kim 2022; Richardson et al. 2021). Some of these semantic features, which we denote as $w_{\text{task}}$, are highly correlated with the original task that the model $\theta_{\text{original}}$ is designed to solve. In addition, for personal identity unlearning tasks, another set of features, $w_{\text{identity}}$, represents information specific to personal information.
@@
(a) Original Model
(b) Error Maximized Model
Figure 2: Feature representations from age classification model. (a) demonstrates clear class distinctions, with age groups well-separated in feature space. (b), derived using Negative Gradient method, shows clustered features with less distinction, illustrating correlation collapse.
Figure 3: The core method of DLFD-feature distribution optimization through optimal transport. This component generates a synthesized dataset by maximizing the distance between retain and forget data distributions in the feature space. When combined with other components (detailed in Algorithm1), DLFD achieves a balance between model utility and forgetting performance.
As illustrated in Figure 2, which visualizes feature representations in the latent space, the Original Model (Figure 2(a)) maintains clear class separations, preserving task-relevant features. In contrast, the Error Maximized model (Figure 2(b)) shows diminished class distinctions, highlighting correlation collapse caused by shared feature disruption.
To address this, we propose Distribution-Level Feature Distancing (DLFD), which maintains task-related features during unlearning (Figure 3). Our method preserves the structure of $\mathcal{W}_{\text{task}}$ while modifying identity-related features.
Proposed Methods
In this section, we introduce Distribution-Level Feature Distancing (DLFD), our comprehensive framework for effective machine unlearning. DLFD consists of three key components designed to balance forgetting performance and model utility.
Feature Distribution Optimization
Our first component shifts the retain data distribution ($\mu$) away from the forget data distribution ($\nu$) by leveraging the optimal transport (OT) distance. Unlike simpler metrics such as KL or JS divergence, it captures the complex, high-dimensional relationships between data points (Arjovsky, Chintala, and Bottou 2017; Gulrajani et al. 2017). The OT distance between the distributions $\mu$ and $\nu$ is defined as:
Here, $\gamma$ is the set of all possible joint distributions that can transport $\mu$ to $\nu$, and $c(w,w^{\prime})$ represents the cost based on cosine similarity between feature vectors. To handle the complexity of solving this problem directly, we employ a differentiable Sinkhorn method (Cuturi 2013; Altschuler, Niles-Weed, and Rigollet 2017), which approximates the solution efficiently and reduces the computational complexity to $\mathcal{O}(n^{2})$ for mini-batch computations.
In this equation, the cost matrix $C$ captures pairwise distances between feature vectors from $\mu$ and $\nu$. The regularization term $\lambda$ keeps the transport plan $T$ smooth, preventing overly concentrated mass transfers that could destabilize the model. Iteratively optimizing $T^{\lambda}$ effectively separates the retain and forget data distributions, mitigating the risk of correlation collapse.
@@
To maintain model utility during the unlearning process, our second component incorporates a classification loss guiding the perturbation process to address correlation collapse. The classification loss ensures that the original class information of the retain data is preserved, even as the model attempts to forget data points. A critical aspect of this component is the use of a linear weight dynamically adjusting the importance of the classification loss throughout the training process.
The linear weight plays a crucial role in balancing the trade-off between maximizing the separation of distributions and preserving the model’s utility. At the beginning of training, it is set lower, allowing the model to focus more on maximizing the distance between the retain and forgot data distributions. As training progresses, the linear weight gradually increases, shifting the model’s focus toward preserving the original class-specific features of the retain data. The perturbation applied to the retain data points $x_{i}$ is computed as follows:
@@
Here, $l_{OT}$ represents the OT loss between the retain and forget data distributions, while $l_{CE}$ indicates the classification loss, weighted by the linear factor $\lambda$ and computed as:
@@
Table 1: Overall performance of various machine unlearning methods on ResNet18 classification tasks. Our method achieves superior NoMUS scores across all tasks, with remarkable forgetting scores while maintaining competitive test accuracy. The best score is in boldface except for the ground-truth (Retrained).
Note: Fine-tuning, NegGrad (Golatkar, Achille, and Soatto 2020b), CF-k, EU-k (Goel, Prabhu, and Kumaraguru 2022a), UNSIR (Tarun et al. 2023), BadTeaching (Chundawat et al. 2023a), SCRUB (Kurmanji, Triantafillou, and Triantafillou 2023)
@@
The goal of an ideal machine unlearning method is to remove the images that need to be forgotten ($\mathcal{D}_{forget}$) while maintaining the original classification performance. In this study, we adopt a common machine unlearning setting where the model has access to a subset of the training data, $\mathcal{D}_{retain}$, which the AI company may still possess. Formally, we assume that the training data $\mathcal{D}_{train}$ is composed of $\mathcal{D}_{retain}$ and $\mathcal{D}_{forget}$, following the general machine unlearning setting described by Choi and Na (2023). Our objective is to develop a machine unlearning algorithm that makes the unlearned model $\theta_{unlearned}$ as similar as possible to the retrained model $\theta_{retrained}$, which is considered the ground truth and is trained only on $\mathcal{D}_{retain}$.
We also introduce a dataset $\mathcal{D}_{unseen}$, which is never used during the training or testing phases of the model. This dataset serves as our test set $\mathcal{D}_{test}$ and is exclusively used for evaluating the forgetting score. It is important to note that any subject targeted for unlearning should not simultaneously belong to the three datasets: $\mathcal{D}_{forget}$, $\mathcal{D}_{retain}$, and $\mathcal{D}_{unseen}$. This ensures that the subject to be forgotten is not present across multiple datasets in the machine unlearning setting.
@@
In this work, we adopt a task-agnostic machine unlearning setup, which ensures that unlearning specific target subjects does not affect the model’s original functionality. Traditional machine unlearning research has primarily focused on class-unlearning, where entire categories (classes) are removed from the model upon a data removal request (Tarun et al. 2023; Golatkar, Achille, and Soatto 2020b; Goel, Prabhu, and Kumaraguru 2022b). While this approach works in certain scenarios, it is not applicable in all cases. For instance, in a gender classification model, removing the male class would leave only the female class, rendering the model ineffective for its intended purpose of gender classification. Hence, class-unlearning is not always representative of real-world needs.
To address these limitations, we propose an instance-unlearning problem setting, which targets the removal of specific personal identities or data samples without changing the overall function of the model. This approach ensures that the model’s core functionality remains intact, making it more applicable to scenarios where the goal is to forget specific data without compromising the model’s utility (Triantafillou et al. 2023; Choi and Na 2023; Choi et al. 2024).
@@
Given $\mathcal{D}_{forget}$ and $\mathcal{D}_{unseen}$ datasets, we train a binary classifier $\psi(\cdot)$ to distinguish between them.
@@
  • Age Estimation: The MUFAC dataset (Choi and Na 2023) contains 13,068 facial images (128$\times$128) in 8 age groups. The training set comprises 10,025 samples, with 8,525 retained and 1,500 designated for forgetting.
  • Emotion Recognition: The RAF-DB dataset (Li, Deng, and Du 2017) contains 15,000 images across 7 emotional classes. The training set comprises 11,044 samples, with 7,730 retained and 3,314 designated for forgetting.
  • Multi-Attribute Classification: The MUCAC dataset (Choi and Na 2023), derived from CelebA (Liu et al. 2018), consists of 30,000 facial images with three binary attributes: gender, age, and expression. The training set includes 25,933 samples, with 15,385 retained and 10,548 for forgetting.
@@
For experiments, we utilize various deep-neural network architectures including ResNet (He et al. 2016), DenseNet (Huang et al. 2017), and EfficientNet (Tan and Le 2019), widely adopted in computer vision. To ensure fair comparison, all machine unlearning methods start from the same $\theta_{original}$ for each task. Specifically, methods fine-tune $\theta_{unlearned}$, initialized as $\theta_{original}$, except for the Retrained model. Serving as ground truth, the Retrained model is trained from scratch on $\mathcal{D}_{retain}$, excluding data to be forgotten, to fully represent the desired unlearning outcome.
Given the computational complexity of our method, which involves calculating OT loss and performing MIA evaluations, we limit the training to a single epoch. Other machine unlearning methods are also trained for 1-2 epochs to ensure a fair comparison. Additionally, we find that learning rates between 0.001 and 0.005 are effective across all models and methods, consistent with previous work (Tarun et al. 2023).
Figure 4: Feature representations from the age classification model trained with DLFD. The model preserves class separation similar to the original model (Figure 2(a)), retaining task-relevant features while mitigating correlation collapse.
@@
Figure 5: The loss distributions for two baselines and ours. The orange space represents the loss distribution for unseen data, while the green represents the loss distribution for forget data. (a) illustrates loss distributions for Original model. (b) shows loss distributions for Retrained model. Finally (c) represents loss distributions for $\theta_{unlearned}$ fine-tuned on DLFD-optimized images.
Performance of DLFD Method
We evaluate our method across four classification tasks: facial age prediction, emotion recognition, multi-attribute classification, and gender classification. The multi-attribute model includes three binary labels: gender (female/male), age (old/young), and expression (smiling/unsmiling), with the average classification accuracy reported as the model utility. Gender classification, originally part of the multi-attribute, is also evaluated as an independent binary classification task.
Across all experiments, our method consistently delivers competitive or superior performance in both metrics. We demonstrate that DLFD effectively unlearns the forget data while maintaining model utility. As shown in Figure 5(c), the loss distributions of $\mathcal{D}{unseen}$ and $\mathcal{D}{forget}$ in our method closely resemble those of the Retrained model, considered the ground truth (Figure 5(b)). This similarity indicates that the unlearning algorithm works effectively.
In particular, the DLFD method shows more substantial improvements in complex, multi-class tasks such as age estimation and emotion recognition, where the feature entanglement is more significant. Conversely, the improvements in multi-attribute and gender classification tasks are relatively smaller, likely due to the binary nature of these classifications, where the complexity of feature entanglement is inherently lower. These results highlight the effectiveness of our approach in scenarios where maintaining feature integrity amid complex and overlapping feature spaces is more challenging.
Moreover, Figure 4 displays the feature representations extracted by the DLFD model. The figure demonstrates how DLFD maintains clear class distinctions, preventing correlation collapse and preserving essential task-related features.
@@
We perform ablation studies to evaluate each component of DLFD. Using only feature distribution optimization initially achieves success in machine unlearning, as shown in Table 3. While effective in separating retain and forget data distributions, this component alone may reduce model utility without the support of other components.
The addition of classification loss preservation as the second component significantly improves performance, especially in Age and Emotion tasks (NoMUS increased by 7.3% and 4.7%), highlighting its role in maintaining model utility. Finally, integrating dynamic forgetting further enhances performance, with additional improvements in Age (2.%) and Emotion (5.3%) tasks. This component effectively prevents correlation collapse by balancing forgetting and utility preservation. The complete framework, combining all three components, achieves superior NoMUS scores across all tasks, showing the effectiveness of their synergistic interaction.
Feature Dist. Optim. Cls Loss Dynamic Forgetting Age Emotion Multi-Attr. Gender NoMUS $\uparrow$ NoMUS $\uparrow$ NoMUS $\uparrow$ NoMUS $\uparrow$ 0.7021 0.7199 0.9423 0.8898 0.7536 0.7536 0.9420 0.9170 0.7698 0.7934 0.9433 0.9192
Table 3: Ablation study results for each component in DLFD, showing cumulative performance improvements.